# Itamar Trainin  Omri Abend

### v2

@@ Abstract
Using LLMs for Multi-Document Topic Extraction has recently gained popularity due to their apparent high-quality outputs, expressiveness, and ease of use. However, most existing evaluation practices are not designed for LLM-generated topics and result in low inter-annotator agreement scores, hindering the reliable use of LLMs for the task. To address this, we introduce $T^{5}Score$, an evaluation methodology that decomposes the quality of a topic set into quantifiable aspects, measurable through easy-to-perform annotation tasks. This framing enables a convenient, manual or automatic, evaluation procedure resulting in a strong inter-annotator agreement score. To substantiate our methodology and claims, we perform extensive experimentation on multiple datasets and report the results. [^1] An implementation of our automatic methodology will be made publicly available upon publication.
@@ 1 Introduction
Still, to substantiate this, LLMs must be assessed carefully. Unfortunately, collecting manually annotated topic sets for direct assessment is impractical (Chang et al., 2009) due to annotators’ bias (Reidsma and op den Akker, 2008; Wich et al., 2020), and the cognitive burden of processing large amounts of information at once (Hoyle et al., 2021; Nugroho et al., 2020). For this reason, existing evaluation procedures rely on indirect approximations (e.g., Chang et al., 2009; Bhatia et al., 2018), rather than direct assessments of quality. Moreover, these practices result in poor Inter-Annotator Agreement (IAA) Stammbach et al. (2023). In the absence of standard evaluation practices, scientific works that employ topic sets often do not include a thorough evaluation of the “correctness” of the extracted topics Livermore et al. (2017); Friese et al. (2018); Keydar et al. (2024). While these practices are common, stipulating the correctness of the topic sets without evidence is unwarranted, and equally in the case of LLM-generated topic sets.
@@ 3.1 Formal Setting & Definitions
$T^{5}Score$ assesses the quality of an FT-topics set by performing 3 annotation tasks (henceforth, measurements): $I(t)$, $R(t,d)$, and $O(t,t^{\prime})$. A measurement is defined as a function of the topic set and the sample and is directly annotated by either a human or a machine (refer to §3.2 for definitions). In this work, we define the explicit functionality of the measurement as annotation guidelines or prompts (refer to Appendix §G, §B).
@@ 3.2 Defining the Quality of an FT-topics Set
According to Friese et al. (2018): “While doing so [theme validation], a constant question needs to be whether the themes … tell a compelling story, or … whether themes need to be collapsed or expanded upon.”. Inspired by this, we say that an FT-topics set is a "good" set concerning a collection of reference documents if it scores high on the following five aspects of quality:
@@ 4.1 USC-SF Survivor Testimonies
A collection of Holocaust survivor testimonies put together by USC-SF. This collection comprises stories recounted by survivors based on their unique experiences and perspectives during the Holocaust. Each testimony naturally describes different experiences, but many of the themes do recur, albeit in a variety of circumstances, times, and places. Therefore, this dataset offers sub-collections of documents on relatively constrained domains, making it ideal for this work. We are further motivated by the use of this dataset in recent computational modeling work (Wagner et al., 2022, 2023; Ifergan et al., 2024; Shizgal et al., 2024).
@@ 5.1 Experimental Setup
We recruited 4 fluent English-speaking in-house annotators. The participants were asked to perform the interpretability ($I(t)$), relevance ($R(t,d)$), and overlap ($O(t,t^{\prime})$) measurements according to the guidelines defined in §3.2 and explicitly described in Appendix G. The annotations were performed over FT-topics sets generated from the USC-SF dataset using GPT3.5 (see Appendix B for the generation prompt). During each session, the annotators were introduced to a set of topics and $M=10$ in-domain reference documents, maintaining full item overlap between annotators. Overall, each annotator was introduced to a total of 21 topic sets and 210 reference documents, see final item counts in Table 1.
@@ F.2 Interpretability Score
Appendix G Annotation Guidelines
@@ 10 EZXckfQkRmxGhcS
Generation Method Rank Generated Desc.
LDA+Prefix W50 5 The theme defined by the following set of words: “int”, “jewish”, “know”, “like”, “jews”, “people”, “went”, “said”, “yes”, “remember”, “mother”, “came”, “us”, “would”, “go”, “jk”, “father”, “well”, “school”, “could”, “fs”, “polish”, “time”, “one”, “non”, “little”, “seconds”, “pauses”, “english”, “think”, “name”, “get”, “yeah”, “used”, “see”, “lot”, “yiddish”, “two”, “war”, “lived”, “never”, “something”, “really”, “home”, “years”, “oh”, “tell”, “say”, “told”, “german”.
@@ 6 Programming and Software Development
Generation Method Rank Generated Desc.
LDA+GPT4 W1 7 Title: Mathematics or Numerical Analysis
@@ 4 Personal Experiences of Discrimination
Generation Method Rank Generated Desc.
GPT4 5 Impact of Nazi Policies

### v3

@@ Abstract
Using LLMs for Multi-Document Topic Extraction has recently gained popularity due to their apparent high-quality outputs, expressiveness, and ease of use. However, most existing evaluation practices are not designed for LLM-generated topics and result in low inter-annotator agreement scores, hindering the reliable use of LLMs for the task. To address this, we introduce $T^{5}Score$, an evaluation methodology that decomposes the quality of a topic set into quantifiable aspects, measurable through easy-to-perform annotation tasks. This framing enables a convenient, manual or automatic, evaluation procedure resulting in a strong inter-annotator agreement score. To substantiate our methodology and claims, we perform extensive experimentation on multiple datasets and report the results. [^1] An implementation of our automatic methodology is available at: https://github.com/itamartrainin/Tpower5Score.
@@ 1 Introduction
Still, to validate such a use, LLMs must be assessed carefully. Unfortunately, collecting manually annotated topic sets for direct assessment is impractical (Chang et al., 2009) due to annotators’ bias (Reidsma and op den Akker, 2008; Wich et al., 2020), and the cognitive burden of processing large amounts of information at once (Hoyle et al., 2021; Nugroho et al., 2020). For this reason, existing evaluation procedures rely on indirect approximations (e.g., Chang et al., 2009; Bhatia et al., 2018), rather than direct assessments of quality. Moreover, these practices result in poor Inter-Annotator Agreement (IAA) Stammbach et al. (2023). In the absence of standard evaluation practices, scientific works that employ topic sets often do not include a thorough evaluation of the “correctness” of the extracted topics Livermore et al. (2017); Friese et al. (2018); Keydar et al. (2024). While these practices are common, stipulating the correctness of the topic sets without evidence is unwarranted, and equally in the case of LLM-generated topic sets.
@@ 3.1 Formal Setting & Definitions
$T^{5}Score$ assesses the quality of an FT-topics set by performing 3 annotation tasks (henceforth, measurements): $I(t)$, $R(t,d)$, and $O(t,t^{\prime})$. A measurement is defined as a function of the topic set and the sample and is directly annotated by either a human or a machine (refer to §3.2 for definitions). In this work, we define the explicit functionality of the measurement as annotation guidelines or prompts (refer to Appendix §H, §B).
@@ 3.2 Defining the Quality of an FT-topics Set
We say that an FT-topics set is a "good" set concerning a collection of reference documents if it scores high on the following five aspects of quality:
@@
3.3 Aggregate Score
While scoring a topic set using individual components is beneficial, in practice, an aggregate score is often preferred. Thus, in Appendix G we present a simple aggregation measure and use it to score the different systems presented in this paper. Nevertheless, as the relative importance of the different components varies considerably between use cases, we advise caution in deciding which aggregation function to use. We defer a deeper discussion of this topic to future work.
@@ 4.1 USC-SF Survivor Testimonies
A collection of Holocaust survivor testimonies put together by USC-SF. [^4] Recorded and transcribed testimonies are available upon request to USC Shoah Foundation. This collection comprises stories recounted by survivors based on their unique experiences and perspectives during the Holocaust. Each testimony naturally describes different experiences, but many of the themes do recur, albeit in a variety of circumstances, times, and places. Therefore, this dataset offers sub-collections of documents on relatively constrained domains, making it ideal for this work. We are further motivated by the use of this dataset in recent computational modeling work (Wagner et al., 2022, 2023; Ifergan et al., 2024; Shizgal et al., 2024).
@@ 5.1 Experimental Setup
We recruited 4 fluent English-speaking in-house annotators. The participants were asked to perform the interpretability ($I(t)$), relevance ($R(t,d)$), and overlap ($O(t,t^{\prime})$) measurements according to the guidelines defined in §3.2 and explicitly described in Appendix H. The annotations were performed over FT-topics sets generated from the USC-SF dataset using GPT3.5 (see Appendix B for the generation prompt). During each session, the annotators were introduced to a set of topics and $M=10$ in-domain reference documents, maintaining full item overlap between annotators. Overall, each annotator was introduced to a total of 21 topic sets and 210 reference documents, see final item counts in Table 1.
@@ F.2 Interpretability Score
Appendix G Aggregate Score
In some use cases, such as quick comparisons between systems or as a reward function for training topic set generation models, an aggregate metric that combines all aspects into a single score is more practical. While the precise formulation may vary across applications, we propose the harmonic mean for the general case due to its simplicity and effectiveness in balancing large and small values, making it well-suited for score averaging. Formally:
$S(\mathcal{T}_{f},\mathcal{D})\text{=}\frac{|A(\mathcal{T}_{f},\mathcal{D})|}{\sum_{\alpha\in A(\mathcal{T}_{f},\mathcal{D})}\frac{1}{\alpha}}$ (9)
where $A(\mathcal{T}_{f},\mathcal{D})$ is the set of aspect scores achieved for the topic set $\mathcal{T}_{f}$ and a set of documents $\mathcal{D}$. Table 9 shows how the different systems fare on the aggregate metric. GPT-4 outperforms all other methods.
Generation Method Dataset
USC SF Multi-News Human Baseline
Random-Letters     0.002 - -
Random-Words     0.001 - -
Domain-Name     $0$
LDA+Prefix W1     0.015 0.114 0.040
LDA+Prefix W10     0.179 0.372 0.186
LDA+Prefix W50     0.150 0.307 0.236
LDA+GPT4 W1     0.008 0.159 0.028
LDA+GPT4 W10     0.098 0.304 0.144
LDA+GPT4 W50     0.158 0.269 0.217
GPT4-Vague     0.134 - -
GPT4-Random     0.149 - -
GPT3.5     0.279 0.435 0.293
GPT4     0.337 0.524 0.612
Table 9: Average aggregate score per generation method and dataset. The best generation method is boldfaced.
Appendix H Annotation Guidelines
@@ 10 EZXckfQkRmxGhcS
5 The theme defined by the following set of words: “int”, “jewish”, “know”, “like”, “jews”, “people”, “went”, “said”, “yes”, “remember”, “mother”, “came”, “us”, “would”, “go”, “jk”, “father”, “well”, “school”, “could”, “fs”, “polish”, “time”, “one”, “non”, “little”, “seconds”, “pauses”, “english”, “think”, “name”, “get”, “yeah”, “used”, “see”, “lot”, “yiddish”, “two”, “war”, “lived”, “never”, “something”, “really”, “home”, “years”, “oh”, “tell”, “say”, “told”, “german”.
@@ 6 Programming and Software Development
7 Title: Mathematics or Numerical Analysis
@@ 4 Personal Experiences of Discrimination
5 Impact of Nazi Policies